Shift summary for DAY shift: 11/11/99

Angela Bellavance (bellavan@rice.edu)
Thu, 11 Nov 1999 15:34:34 -0600 (CST)

Shift Summary for day shift, 11/Nov/99

On Shift: Leo Bellantoni, Angela Bellavance

Hours of beam delivered at all intensities: 0
Hours of beam delivered at nominal intensity: 0
Hours of E799 data (special runs not included): 0
Hours of special runs (laser scans, muon runs...): 0.5

Accounting of lost time: TeV down for studies

========================================================================
Runs ended this shift: Run 14942 - LASER SCAN

Tapes brought to FCC this shift: none

========================================================================
Accesses during shift:

Jason LaDue to pick up some tools.

Bud to check on which shield blocks to remove for installing the KAMI
calorimeter. He notices one of the warning lights on the magnet is burned
out.

Fire alarm people to clean and test the alarms.

Hogan and Ashkan to put the DC4 low voltage power supply on remote
control.

Doug J. to take laser alignment readings and fix a hissing vacuum valve.

Bob Hsiung to pick up 25 PMT bases.

Sasha Golossanov and Valeri to bring in equipment for Aerogel testing.

Jason LaDue and Hogan to fix DC4 readout problem(s).

Angela to replace light bulb on analysis magnet.

========================================================================
Overview of shift:
Hall is abuzz with people doing tasks on BobT's list - trying to get
detector running well again for return of beam.

========================================================================
Description of Shift:

time comment
------------------------------------------------------------------------
8:06 Analysis magnet turned off per request by people working
on the interlock. Run initing is hanging in level 3. Experts
working on it.
9:11 Begin Laser Scan (finishes at 9:43).
12:49 Shift checklist modified to include running the kumac "montup799".
14:00 Fire alarm people trip off CsI low voltage. Leo turns it back on
with help from Julie.
14:45 Angela does her best golden retriever impression and fetches a
light bulb for the magnet from the Proton Pole Building.
(Extra good bulb placed in large red tool cabinet upstairs,
phone/electrical drawer.)
15:20 Gas checklist done.

========================================================================
Follow up Notes on Previous Shift:

Laser scan done!

Coffee machine fixed!

========================================================================
Items to Follow up on Next Shift:

Don't forget to run "montup799" and do everything else on the shift
checklist!

How many KTeVers does it take to change a light bulb?
.

.

.
Two! Everyone knows you can't go on controlled access by yourself!!

_______
Angela Bellavance work: (630) 840-4144
bellavan@rice.edu home: (630) 548-3412
http://www.owlnet.rice.edu/~bellavan/ page: (630) 722-6861
"It has yet to be proven that intelligence has any survival value."
- Arthur C. Clarke